OpenStudy (anonymous):

14y^2-686/4y^2-196= 7/2? 14y^2-686/4y^2-196= 7/2? @Mathematics

OpenStudy (anonymous):

again try to simplify the terms

OpenStudy (anonymous):

\[\frac{14y^2-686}{4y^2-196} = \frac{7}{2}\]

OpenStudy (anonymous):

thats what i have is 7/2

OpenStudy (anonymous):

fram numerator take 14 common

OpenStudy (anonymous):

.......i was wrong

OpenStudy (anonymous):

yes love u were wronf

OpenStudy (anonymous):

thats not right?

OpenStudy (anonymous):

still no solution

OpenStudy (anonymous):

(14y^2-686)/(4y^2-196)=1

OpenStudy (anonymous):

hmm idk i keep getting 7/2

OpenStudy (anonymous):

here we will get \[\frac{1}{4} = \frac{1}{4}\]

OpenStudy (anonymous):

sheg, simplify the left side of the equation, you will get 1

OpenStudy (anonymous):

tena, 7/2 is your answer? if it is your answer, then the answer will be 0